In a company having 80 employees, 60 earn 10.00 per hour and 20 earn $13.00 per hour. Determine the mean earning per hour.b would the answer in part (a) be the same if the 60 employees earn a mean hourly wage.


1
Expert's answer
2022-05-30T23:03:38-0400

a)


"mean=\dfrac{n_1x+n_2y}{n_1+n_2}=\dfrac{60(10)+20(13)}{60+20}=10.75"

b)

If "n_1" employees earn a mean hourly wage "\bar{x}" then their total amount earned is "n_1\bar{x}."

If "n_2" employees earn "y" per hour then their total amount earned is "n_2y."

Then


"mean=\dfrac{n_1\bar{x}+n_2y}{n_1+n_2}=\dfrac{60(10)+20(13)}{60+20}=10.75"

The answer would be the same.